A 1.5dm3 of potassium chloride (KCl) has a concentration of 0.5mol/dm3.
How many grams of potassium chloride are present?
Given the atomic masses:
K = 39 and Cl = 35.5
1
Expert's answer
2018-09-06T07:27:04-0400
n = C*V
n(KCl) = 0.5mol/dm3*1.5dm3 = 0.75 mol
M(KCl) = 39 + 35.5 = 74.5 g/mol
m(KCl) = n*M = 0.75mol*74.5g/mol = 55.875 g